# Simulating toponium formation signals at the LHC

### v1

@@ 1 Introduction
The remainder of this study is organised as follows. In section 2, we outline how to modify matrix elements relevant to the production of a colour-singlet top-antitop quark pair in hadronic collisions to account for toponium contributions through Green’s function ratios. Section 3 details the implementation of these modifications in the MG5_aMC event generator Alwall:2014hca , enabling fixed-order predictions matched with parton showers to supplement conventional Standard Model (SM) top-antitop predictions, which typically neglect non-perturbative contributions. In sections 4 and 5, we apply this strategy to study a few distributions highlighting the impact of the toponium effects, first at the level of hard-scattering events and next at LO+PS. Finally, in section 6, we summarise our findings and provide an outlook on further improvements to enhance the accuracy of SM top-antitop production predictions.
@@ 2 Non-relativistic QCD Green’s function for toponium formation
As an example, in our previous work Fuks:2021xje we adopted a Coulombic potential including a running strong coupling constant normalised from its value at the $Z$-pole. This choice for the potential provides a good starting point for studying toponium formation, as it allows for analytical control of the results by freezing the value of the coupling constant at a specific scale. Additionally, it serves as a useful cross check of our numerical code, which is general and can handle any form of the QCD potential $\widetilde{V}_{\mathrm{QCD}}(\vec{p})$ in the momentum space. We determined the Green’s function $\widetilde{G}(E;p)$ for the QCD potential (12) with a running strong coupling constant normalised at $\alpha_{s}(m_{Z})=0.12$, and a top quark mass $m_{t}=173$ GeV and width $\Gamma_{t}=1.49$ GeV. [^1] In order to avoid numerical issues caused by the infrared pole of the QCD running coupling, we applied a Richardson-like regularisation Richardson:1978bt , following the methodology outlined in Jezabek:1992np . Thanks to the physical infrared cutoff provided by the top quark width $\Gamma_{t}\approx 1.5$ GeV, the results show minimal sensitivity to the long distance behaviour of the QCD potential. The results are given in two data files, swData.M20.M5.txt and swData.M5.P20.txt, which contain tabulated information on ratios of the Green’s functions $\widetilde{G}(E;p)/\widetilde{G}_{0}(E;p)$. This information is provided in the form of two-dimensional grids in $(E,p)$. The first file focuses on low binding energy values, containing 150 points for $E$ in the range $-20$ GeV $\leq E\leq-5$ GeV, and 400 points for $p$ in the range $0<p\leq 100$ GeV. The second file addresses larger binding energy values, with 1250 points for $E$ in the range $-5$ GeV $\leq E\leq 20$ GeV, and the same 400 points for $p$. Each line in the files corresponds to specific $E$ and $p$ values, and contains four entries, the binding energy $E$, the momentum $p$ and the real and imaginary part of the ratio of the Green’s functions. These files are available from a public repository at https://github.com/BFuks/toponium.git.
Figure 1: Norm of the Green’s function $|\widetilde{G}(E;p)|^{2}$ in the presence of the Coulombic potential (12) as obtained after fixing the strong coupling constant at the $Z$-pole $\alpha_{s}(m_{Z})=0.12$ (left), and norm of the free Green’s function (11), $|\widetilde{G}_{0}(E;p)|^{2}$ (right). The results are shown as a function of the binding energy $E$ and the momentum $p$.
@@
for $m_{t}=173$ GeV. This behaviour is approximately recovered for the interacting Green’s function at positive binding energies ($E\gtrsim 2$ GeV), though with a different overall normalisation that persists until significantly larger values of $E$ (not shown on the figure). On the other hand, for $p\lesssim 20$ GeV and $-4$ GeV $\lesssim E\lesssim 1$ GeV, the toponium wave function not only enhances the magnitude of the Green’s function squared norm, but also substantially distorts its shape, with the most prominent effect occurring around $E\simeq 2$ GeV at low momentum $p$.
@@ 5 Toponium phenomenology at the LHC
In this section, we employ the simulation chain introduced in section 3 to generate 1,000,000 toponium events at LO+PS, either incorporating the re-weighting (15) or not, together with the projection of the matrix element into a colour-singlet state. We use MG5_aMC Alwall:2014hca for hard-scattering simulations with the CT18NLO set of parton densities Hou:2019efy ; Buckley:2014ana , and Pythia 8 Bierlich:2022pfr for parton showering and hadronisation. The reconstruction of events is performed using the anti-$k_{T}$ algorithm Cacciari:2008gp with a jet radius parameter of $R=0.4$, as implemented in FastJet Cacciari:2011ma and interfaced with MadAnalysis 5 Conte:2018vmg ; Araz:2020lnp . The latter software is also employed to apply the selection strategy previously introduced in Fuks:2021xje . We then compare our predictions for the toponium signal with the background from conventional top-antitop pair production, followed by di-leptonic decays. We have simulated 3,100,000 background events at LO+PS accuracy, with a total cross section normalised to next-to-next-to-leading order with next-to-next-to-leading logarithmic threshold resummation (NNLO+NNLL) Czakon:2013goa , using thus a cross section value of $\sigma_{t\bar{t}}=36.9$ pb.
@@
Figure 6: Distributions of the di-lepton azimuthal angle difference $\Delta\varphi_{\bar{\ell}\ell^{\prime}}$ (left) and pseudo-rapidity difference $\Delta\eta_{\ell\ell^{\prime}}$ (right), emerging from di-leptonic $t\bar{t}$ production and decay at a centre-of-mass energy $\sqrt{s}=13$ TeV with an integrated luminosity of 140 fb-1. Predictions are shown for conventional top-antitop production without toponium effects (orange pentagons) and for toponium production with Green’s function re-weighting (teal hexagons).
Figure 7: Same as in figure 6, but for the di-lepton separation in the transverse plane, $R_{\bar{\ell}\ell^{\prime}}$.
@@
Toponium characterisation can be further achieved from a kinematic reconstruction of the system made of the two leptons, $b$-jets and missing transverse momentum, as detailed in our previous study Fuks:2021xje . This would require pairing each lepton $\ell_{1,2}$ with a $b$-jet. Here, $\ell_{1}$ and $\ell_{2}$ refer to the hardest and softest leptons, respectively, while we denote by $b_{i}$ the $b$-jet paired with lepton $\ell_{i}$. The correct pairing is determined by requiring that $m(\ell_{1},b_{1})>m(\ell_{1},b_{2})$. Assuming that the transverse momenta of the reconstructed top and antitop are approximately equal, which should hold for the signal, we can proceed to reconstruct the four-momenta of the two neutrinos through a kinematic fit. The resulting second-order equation for each neutrino introduces a two-fold ambiguity, which we resolve by choosing the solution that yields a reconstructed top mass closest to the known $m_{t}$ value. Examining distributions of observables involving the two reconstructed top quarks could then potentially be used as handles to characterise the excess events observed by the ATLAS and CMS collaborations ATLAS:2019hau ; ATLAS:2023gsl ; CMS:2024ybg as arising from toponium. We refer to Fuks:2021xje for more details.
@@ Acknowledgements.
We thank Yuki Sumino for fruitful discussions on kinematic distributions of toponium-induced final states. The work of BF was supported in part by Grant ANR-21-CE31-0013 (Project DMwithLLPatLHC) from the French Agence Nationale de la Recherche (ANR), the one of KM by a Natural Science Basic Research Program of Shaanxi Province, China (2023-JC-YB-041), and the one of YJZ by JSPS KAKENHI Grant No. 21H01077 and 23K03403.

### v2

@@ 1 Introduction
The remainder of this study is organised as follows. In section 2, we outline how to modify matrix elements relevant to the production of a colour-singlet top-antitop quark pair in hadronic collisions to account for toponium contributions through Green’s function ratios. Section 3 details the implementation of these modifications in the MG5_aMC event generator Alwall:2014hca , enabling fixed-order predictions matched with parton showers to supplement conventional SM top-antitop predictions, which typically neglect non-perturbative contributions. In sections 4 and 5, we apply this strategy to study a few distributions highlighting the impact of the toponium effects in LHC collisions at a centre-of-mass energy of $\sqrt{s}=13$ TeV, first at the level of hard-scattering events and next at LO+PS. Finally, in section 6, we summarise our findings and provide an outlook on further improvements to enhance the accuracy of SM top-antitop production predictions.
@@ 2 Non-relativistic QCD Green’s function for toponium formation
As an example, in our previous work Fuks:2021xje we adopted a tree-level Coulombic potential including a strong coupling constant normalised from its value at the $Z$-pole. This choice for the potential provides a good starting point for studying toponium formation, as it allows for analytical control of the results by freezing the value of the coupling constant at a specific scale. Additionally, it serves as a useful cross check of our numerical code, which is general and can handle any form of the QCD potential $\widetilde{V}_{\mathrm{QCD}}(\vec{p})$ in the momentum space. We determined the Green’s function $\widetilde{G}(E;p)$ for the tree-level part of the QCD potential (12) with a running strong coupling constant normalised at $\alpha_{s}(m_{Z})=0.12$, and a top quark mass $m_{t}=173$ GeV and width $\Gamma_{t}=1.49$ GeV. [^1] In order to avoid numerical issues caused by the infrared pole of the QCD running coupling, we applied a Richardson-like regularisation Richardson:1978bt , following the methodology outlined in Jezabek:1992np . Thanks to the physical infrared cutoff provided by the top quark width $\Gamma_{t}\approx 1.5$ GeV, the results show minimal sensitivity to the long distance behaviour of the QCD potential. The results are given in two data files, swData.M20.M5.txt and swData.M5.P20.txt, which contain tabulated information on ratios of the Green’s functions $\widetilde{G}(E;p)/\widetilde{G}_{0}(E;p)$. This information is provided in the form of two-dimensional grids in $(E,p)$. The first file focuses on low binding energy values, containing 150 points for $E$ in the range $-20$ GeV $\leq E\leq-5$ GeV, and 400 points for $p$ in the range $0<p\leq 100$ GeV. The second file addresses larger binding energy values, with 1250 points for $E$ in the range $-5$ GeV $\leq E\leq 20$ GeV, and the same 400 points for $p$. Each line in the files corresponds to specific $E$ and $p$ values, and contains four entries, the binding energy $E$, the momentum $p$ and the real and imaginary part of the ratio of the Green’s functions. These files are available from a public repository at https://github.com/BFuks/toponium.git.
Figure 1: Norm of the Green’s function $|\widetilde{G}(E;p)|^{2}$ in the presence of the tree-level Coulombic potential as obtained after fixing the strong coupling constant at the $Z$-pole $\alpha_{s}(m_{Z})=0.12$ (left), and norm of the free Green’s function (11), $|\widetilde{G}_{0}(E;p)|^{2}$ (right). The results are shown as a function of the binding energy $E$ and the momentum $p$.
@@
for $m_{t}=173$ GeV. This behaviour is approximately recovered for the interacting Green’s function at positive binding energies ($E\gtrsim 2$ GeV), though with a different overall normalisation that persists until significantly larger values of $E$ (not shown on the figure). On the other hand, for $p\lesssim 20$ GeV and $-4$ GeV $\lesssim E\lesssim 1$ GeV, the toponium wave function not only enhances the magnitude of the Green’s function squared norm, but also substantially distorts its shape, with the most prominent effect occurring around $E\simeq-2$ GeV at low momentum $p$.
@@ 5 Toponium phenomenology at the LHC
In this section, we employ the simulation chain introduced in section 3 to generate 1,000,000 toponium events at LO+PS for proton-proton collisions at a centre-of-mass energy of $\sqrt{s}=13$ TeV, either incorporating the re-weighting (15) or not, together with the projection of the matrix element into a colour-singlet state. We use MG5_aMC Alwall:2014hca for hard-scattering simulations with the CT18NLO set of parton densities Hou:2019efy ; Buckley:2014ana , and Pythia 8 Bierlich:2022pfr for parton showering and hadronisation. The reconstruction of events is performed using the anti-$k_{T}$ algorithm Cacciari:2008gp with a jet radius parameter of $R=0.4$, as implemented in FastJet Cacciari:2011ma and interfaced with MadAnalysis 5 Conte:2018vmg ; Araz:2020lnp . The latter software is also employed to apply the selection strategy previously introduced in Fuks:2021xje . We then compare our predictions for the toponium signal with the background from conventional top-antitop pair production, followed by di-leptonic decays. We have simulated 3,100,000 background events at LO+PS accuracy, with a total cross section normalised to next-to-next-to-leading order with next-to-next-to-leading logarithmic threshold resummation (NNLO+NNLL) Czakon:2013goa , using thus a total cross section value of $\sigma_{t\bar{t}}=36.9$ pb for the 2-to-6 scattering process.
@@
Figure 6: Distributions of the di-lepton azimuthal angle difference $\Delta\varphi_{\ell\ell^{\prime}}$ (left) and pseudo-rapidity difference $\Delta\eta_{\ell\ell^{\prime}}$ (right), emerging from di-leptonic $t\bar{t}$ production and decay at a centre-of-mass energy $\sqrt{s}=13$ TeV with an integrated luminosity of 140 fb-1. Predictions are shown for conventional top-antitop production without toponium effects (orange pentagons) and for toponium production with Green’s function re-weighting (teal hexagons).
Figure 7: Same as in figure 6, but for the di-lepton separation in the transverse plane, $R_{\ell\ell^{\prime}}$.
@@
Toponium characterisation can be further achieved from a kinematic reconstruction of the system made of the two leptons, $b$-jets and missing transverse momentum, as detailed in our previous study Fuks:2021xje . This would require pairing each lepton $\ell_{1,2}$ with a $b$-jet. Here, $\ell_{1}$ and $\ell_{2}$ refer to the hardest and softest leptons, respectively, while we denote by $b_{i}$ the $b$-jet paired with lepton $\ell_{i}$. The correct pairing is determined by requiring that $m(\ell_{1},b_{1})>m(\ell_{1},b_{2})$. Assuming that the transverse momenta of the reconstructed top and antitop are approximately equal, which should hold for the signal, we can proceed to reconstruct the four-momenta of the two neutrinos through a kinematic fit. The resulting second-order equation for each neutrino introduces a two-fold ambiguity, which we resolve by choosing the solution that yields a reconstructed top mass closest to the known $m_{t}$ value. Examining distributions of observables involving the two reconstructed top quarks could then potentially be used as handles to characterise the excess events observed by the ATLAS and CMS collaborations ATLAS:2019hau ; ATLAS:2023gsl ; CMS:2024ybg as arising from toponium. We refer to Fuks:2021xje ; Aguilar-Saavedra:2024mnm for more details.
@@ Acknowledgements.
We thank Yuki Sumino for fruitful discussions on kinematic distributions of toponium-induced final states, and Thomas Waluda-Duquesne for useful comments on our Green’s function tables. The work of BF was supported in part by Grant ANR-21-CE31-0013 (Project DMwithLLPatLHC) from the French Agence Nationale de la Recherche (ANR), the one of KM by a Natural Science Basic Research Program of Shaanxi Province, China (2023-JC-YB-041), and the one of YJZ by JSPS KAKENHI Grant No. 21H01077 and 23K03403.